# An Empirical Study of Self-Admitted Technical Debt in Machine Learning Software

### v2

@@ 1. Introduction
The burgeoning of the Machine Learning (ML) era signals increased deployment of ML software in a wide range of software domains including large scale systems for healthcare, mobility, banking, transportation, policing, marketing, and social media (jordan2015machine, ). ML-based systems, i.e., projects that integrate ML models with other non-ML components, serve as backbone for this expansive integration across domains.
@@
Unless they reuse third-party models, ML-based systems involve the training of custom ML models, which is achieved using a ML pipeline composed of multiple stages like data preprocessing, ML model training, model evaluation, and deployment (bhatia2023towards, ). The development of ML-based systems necessitates continuous management of ML-related assets (idowu2021asset, ) (e.g., training data, ML models, job execution data) and introduces additional challenges that are not prevalent in traditional (non-ML) software (sato2019continuous, ). While popular ML toolkits such as Scikit-Learn, TensorFlow, and Auto ML, have enabled developers to easily leverage state-of-the-art ML algorithms, the specification of ML algorithms is far from trivial (d2022underspecification, ), and development and maintenance practices for software built around ML differ from those of traditional software products (wan2019does, ). For instance, the dependency of ML code on external data makes it difficult to enforce strict abstraction boundaries (Sculley, ). Apart from ML toolkit frameworks, ML-based systems also need (amongst others) established protocols for iterative model training, maintaining a structured log of experiments and their outcomes, while pruning dead code paths to promptly maintain codebase clarity.
@@
As a result of these challenges and the tight deadlines for rushing AI products to the market, developers and data scientists can make suboptimal decisions for short-term benefits (obrien202223, ) when working on their code, models and data, a phenomenon described by Cunningham (cunningham1992wycash, ) as technical debt (TD). In software engineering, TD manifests as 1) optimization issues where optimization for short-term puts long-term into technical (and economic) jeopardy, 2) short-cuts in software implementation giving the perception of success until their consequences have crippling effects, and 3) development and architectural decisions prioritizing short-term causing long-term consequences (brown2010managing, ). TD may represent inferior development implementations in software design, missing/incomplete documentation, known defects, code, testing, or software requirements as identified as Bavota et al. (bavota2016large, ) for traditional software; or as data dependencies, hidden feedback loops or pipeline jungles for ML software as identified by Sculley et al. (Sculley, ). TD results from decisions made while evaluating competing development tradeoffs; these decisions, just like financial debt, will require correction in the future (klinger2011enterprise, ). Such a correction may come with non-trivial interest, for instance, as identified by Guo et al. (guo2016exploring, ), it can take up to 100% interest probability, i.e., taking as many hours to pay off the TD as the principal time taken to develop the initial software artifact.
@@
TD in general is hard to automatically detect from software engineering resources. However, in an important subset of cases, developers explicitly admit having produced a somewhat temporary implementation that possibly requires more work in the future. Such debt is referred to as Self-Admitted Technical Debt (SATD), and can be identified within comments that are inserted in the source code while developing (sierra2019survey, ). For example, Figure 1 provides an example of a natural language processing application, Mead-ML [^1] Accessed on: 2024-03-22, https://github.com/mead-ml/mead-baseline/commit/d4d8fab88a7424785ef508d0c1bd24ac58cae4d8#diff-c8db4ed3ecd32411a0c3bfecf8555d4ca0cf2181bfcce78b670b9cdc79e0b99d, where the code has a “dirty” implementation of saving a dummy matrix of zeros. In lines #190 and #192, the developers admit that this ad-hoc implementation (TD) will be fully developed after “check pointing” is implemented in the future. Another example of SATD declaration is the comment “# TODO handle attention history [^2] Accessed on: 2024-03-22, https://github.com/ufal/neuralmonkey/blob/master/neuralmonkey/attention/transformer_cross_layer.py within a code base that builds an attention model in the Neural Monkey project, since it indicates the missing implementation of handling model weights.
@@
The problem of TD in ML is exacerbated due to the “implicit” requirements of ML software products, which are grounded in data rather than specification documentations. As pointed out by Sculley et al. (Sculley, ), the problem of TD is pronounced in ML due to the stochastic nature of ML, as opposed to the deterministic nature of traditional (non-ML) software, leading to quality issues similar to traditional, non-ML software as well as new types of issues. For instance, an SATD in an ML model’s hyper-parameter configurations or data pre-processing will result in an inferior model with lower accuracy. Every incorrect decision made by such a model may potentially negatively affect various aspects such as the vendor’s reputation, user experience, or financial performance. Building on the insights of Sculley et al. (Sculley, ), this paper sets out to empirically validate whether the technical debt manifestations they describe in earlier work are indeed prevalent in real-world ML applications-and if so what are their characteristics. Empirically understanding the nature of SATD via code comments as recommended by Potdar and Shihbab allows us to explore direct insights into the acknowledged shortcomings in the implementation of ML systems. Since the nature of code changes performed in ML pipelines is different than that of the traditional non-ML software (bhatia2023towards, ), it is only reasonable to expect that SATD in ML code should appear distinct. Hence, we use code comments as a starting point, whereas other aspects of ML technical debt, such as debt in the underlying data, or model artifacts (perhaps embedded in project documentation, architectural design documentation, or project management systems) fall outside the scope of this current study. This delineation lays the groundwork for future research to explore such avenues.
@@
Despite recent efforts to understand SATD introduction and removal in deep learning frameworks, we still do not know much about the diffusion and evolution of technical debt in ML-based systems using such frameworks. Neither do we know the composition or the circumstances of SATD introduction or removal. While several researchers have investigated the detection, evolution, and impact of technical debt in traditional software systems (e.g., (Alfayez2020, ; Lim2012, ; liu2021exploratory, ; wehaibi2016examining, ; brown2010managing, ; letouzey2012managing, ; kruchten2012technical, ; kruchten2013technical, ; bavota2016large, )), software systems powered by ML models (i.e., ML-based systems) have mostly remained out of scope.
@@
Figure 1. Example of SATD(s) in a Natural Language Processing application, Mead-ML.
@@ 2.1. Studies on Software Engineering for Machine Learning
A systematic literature review by Nascimento et al. (nascimento2020software, ) highlighted the unique challenges in ML testing, ML software quality, and data management in ML systems. Moin et al. (moin2021data, ) proposed a novel approach to integrate Model-Driven Software Engineering and Model-Driven ML Engineering, with a focus on automated ML, aiming to assist software engineers without deep ML knowledge. Tantithamthavorn et al. (tantithamthavorn2018impact, ) emphasized the need for explainable ML in software engineering, making ML models more practical and actionable. Kästner and Kang (kastner2020teaching, ) discussed the importance of teaching software engineering skills to students with an ML background, focusing on realism, robust infrastructure, and ethical considerations. Singla et al. (singla2020story, ) analyzed the adoption of Agile methodology in machine learning projects, finding a higher number of exploratory tasks and difficulty in estimating task duration.
Within such a vast body of SE4ML research, our research expands on the “deployment maintenance and life-cycle management” area. This area is a ripe domain with prior research focusing on various aspects like code changes (e.g., Bhatia et al. (bhatia2023towards, ) examined the nature of code changes in open source ML software), ML asset management (Idowu et al. (idowu2022asset, ) focused on tracking, exploring, and retrieving ML assets), ML code smells (e.g., Jebnoun et al. (jebnoun2020scent, ) analyzed 10 code smells and their properties in ML software). In contrast, our study focuses on the SATD aspect and helps ML stakeholders understand the nature of TD and enable ML practitioners to gain awareness of TD in building ML applications.
@@ 2.2. Empirical Studies on SATD in Traditional (Non-ML) Software
In 1994, Cunningham (cunningham1992wycash, ) introduced the concept of technical debt as “quick and dirty” work in software design and implementation resulting in “not quite-right code”. However, TD is an abstract concept, as making tradeoffs between optimal software and meeting project deadlines is not quantitatively measurable. Hence, subsequently, Potdar and Shibab (potdar2014exploratory, ) extended the concept of TD to self-admitted TD (SATD), which was deemed to be a more visible measure of TD (nord2012search, ; kruchten2012technical, ; Alves2016, ). Along with introducing the concept of SATD, Potdar and Shibab examined SATDs in large open-source projects - Eclipse, Chromium, Apache HTTP, and ArgoUML. They found a high prevalence (within up to 31% of files) of SATD, and identified that more experienced developers add more SATD and are unable to pay off the technical development due to time pressure and code complexity.
@@
Building on the seminal work of Potdar and Shibab, multiple subsequent studies have identified SATD in traditional software. For instance, Maldonado et al. (maldonado2017empirical, ) studied the evolutionary aspects of traditional code and revealed that SATD lasts between 18 and 172 days in a system. Bavota and Russo (bavota2016large, ) performed a large-scale study of SATD across 159 software projects, contributing towards developing a taxonomy of SATD for traditional software. Wehaibi et al. (wehaibi2016examining, ) examined the impact of SATD on software quality, while Kashiwa et al. (kashiwa2022empirical, ) studied SATD in code reviews.
@@
However, none of the prior studies investigate SATD in ML-based systems. As identified by the previous subsection, the engineering practices and research-based findings of traditional software cannot be expected to hold as-is for ML-based systems (washizaki2019studying, ; ozkaya2020really, ; martinez2021software, ), and hence, our research is important to identify the characteristics of SATD in the currently prevalent ML-era of software development.
@@
Two studies bear similarities to our research. Firstly, akin to our study, the study by Liu et al. (liu2018satd, ) explores SATD in deep learning frameworks by investigating the introduction and removal of SATD in seven deep learning frameworks. However, contrary to the study of frameworks, our research identifies debt in ML systems, which use the former frameworks to build and integrate pipelines into ML applications. Our study provides insights for a broader spectrum of ML stakeholders who do not deal with ML mathematical or algorithmic implementations, but need to manage ML assets (like data, model) for development of ML applications. Secondly, the study by O’Brien et al. (obrien202223, ) mirrors our work closely. The authors check the different types of SATDs in ML software and check their distribution in different components of ML pipeline stages. While O’Brien et al.’s research is qualitative, we focus on both quantitative and qualitative aspects of SATD on a different dataset. Within the quantitative aspects, our work takes a genealogical approach to SATD, tracing the lifecycle from introduction to removal. Additionally, we put forth an exploratory model to understand the dynamics of long-lasting TD. Given our different dataset, our qualitative results complement the results of O’Brien et al. and provide a broader understanding of the prevalence and impact of debt across different datasets and domains.
@@ 2.3. Studies Presenting SATD Detection Tools
A few notable examples are the efforts made by Malando et al. (da2017using, ), who constructed an NLP-based SATD classification model and tested their tool across ten substantial open-source projects. Later, in 2017, Huang et al. (huang2018identifying, ) expanded the scope by developing a predictive SATD detection ML model that utilizes feature selection to provide the important words from the comment text used for SATD identification.
As an outcome of this research, the authors presented a composite SATD identification tool that combines multiple classifiers from different source code projects. Subsequently, Ren et al. (ren2019neural, ) utilized neural networks for SATD identification by leveraging the Convolutional Neural Network architecture to deliver state-of-the-art performance in detecting SATD comments. Later, in 2020, Faris et al. (de2020identifying, ) enhanced SATD classification models by using a contextualized vocabulary that considers the level of importance of different textual patterns along with the relationship between patterns and debt types.
@@
We employed the text mining-based SATD detector tool developed by Liu et al. (liu2018satd, ) in our study, which leverages natural language processing and machine learning to classify SATD comments. We selected Liu et al.’s SATD detector tool because of its frequent maintenance and demonstrated applicability in other authoritative SATD studies (e.g.,  (liu2021exploratory, ) and (Zampetti, )).
@@ 3.1. RQ1: What is the prevalence of SATD in ML-based systems?
Sculley et al.’s seminal work (Sculley, ) highlighted the tendency of ML programs to incur technical debt, remarking that the strong dependence of ML models on data makes it difficult to enforce strict abstraction boundaries. When building ML-based systems, developers often experiment with different configurations in search of the optimal model; building, testing, and comparing the performance of different prototypes to identify the best configuration leading to the most efficient model (amershi2019software, ). This process can often result in multiple dead experimental code paths, increasing the cyclomatic complexity of the code and the risk of an unexpected behavior arising from some of the obsolete experimental code paths (Sculley, ). Another potential area where technical debt can accumulate quickly in ML code is in the model building or data processing configurations. Developers often have to handle a wide range of configuration options. For example, they often have to select among a variety of algorithm-specific learning settings, and–or from a large pool of pre-processing, post-processing, and verification techniques. In this research question, we aim to examine the extent to which ML software deployed today, contains technical debt, by comparing the prevalence of SATD in ML and non-ML code. While Sculley et al. (Sculley, ) elucidated additional ML challenges leading to accumulation of technical debt that are not present in non-ML software, empirical evaluations of said challenges still remain scarce. Hence, in this research question, we aim to examine the extent to which ML software deployed today, contains technical debt, by comparing the prevalence of SATD in ML and non-ML code.

### v3

@@ 1. Introduction
The burgeoning of the Machine Learning (ML) era signals increased deployment of ML software in a wide range of software domains including large scale systems for healthcare, mobility, banking, transportation, policing, marketing, and social media (jordan2015machine). ML-based systems, i.e., projects that integrate ML models with other non-ML components, serve as backbone for this expansive integration across domains.
@@
Unless they reuse third-party models, ML-based systems involve the training of custom ML models, which is achieved using a ML pipeline composed of multiple stages like data preprocessing, ML model training, model evaluation, and deployment (bhatia2023towards). The development of ML-based systems necessitates continuous management of ML-related assets (idowu2021asset) (e.g., training data, ML models, job execution data) and introduces additional challenges that are not prevalent in traditional (non-ML) software (sato2019continuous). While popular ML toolkits such as Scikit-Learn, TensorFlow, and Auto ML, have enabled developers to easily leverage state-of-the-art ML algorithms, the specification of ML algorithms is far from trivial (d2022underspecification), and development and maintenance practices for software built around ML differ from those of traditional software products (wan2019does). For instance, the dependency of ML code on external data makes it difficult to enforce strict abstraction boundaries (Sculley). Apart from ML toolkit frameworks, ML-based systems also need (amongst others) established protocols for iterative model training, maintaining a structured log of experiments and their outcomes, while pruning dead code paths to promptly maintain codebase clarity.
@@
As a result of these challenges and the tight deadlines for rushing AI products to the market, developers and data scientists can make suboptimal decisions for short-term benefits (obrien202223) when working on their code, models and data, a phenomenon described by Cunningham (cunningham1992wycash) as technical debt (TD). In software engineering, TD manifests as 1) optimization issues where optimization for short-term puts long-term into technical (and economic) jeopardy, 2) short-cuts in software implementation giving the perception of success until their consequences have crippling effects, and 3) development and architectural decisions prioritizing short-term causing long-term consequences (brown2010managing). TD may represent inferior development implementations in software design, missing/incomplete documentation, known defects, code, testing, or software requirements as identified as Bavota et al. (bavota2016large) for traditional software; or as data dependencies, hidden feedback loops or pipeline jungles for ML software as identified by Sculley et al. (Sculley). TD results from decisions made while evaluating competing development tradeoffs; these decisions, just like financial debt, will require correction in the future (klinger2011enterprise). Such a correction may come with non-trivial interest, for instance, as identified by Guo et al. (guo2016exploring), it can take up to 100% interest probability, i.e., taking as many hours to pay off the TD as the principal time taken to develop the initial software artifact.
@@
TD in general is hard to automatically detect from software engineering resources. However, in an important subset of cases, developers explicitly admit having produced a somewhat temporary implementation that possibly requires more work in the future. Such debt is referred to as Self-Admitted Technical Debt (SATD), and can be identified within comments that are inserted in the source code while developing (sierra2019survey). For example, Figure 1 provides an example of a natural language processing application, Mead-ML [^1] Accessed on: 2024-03-22, https://github.com/mead-ml/mead-baseline/commit/d4d8fab88a7424785ef508d0c1bd24ac58cae4d8#diff-c8db4ed3ecd32411a0c3bfecf8555d4ca0cf2181bfcce78b670b9cdc79e0b99d, where the code has a “dirty” implementation of saving a dummy matrix of zeros. In lines #190 and #192, the developers admit that this ad-hoc implementation (TD) will be fully developed after “check pointing” is implemented in the future. Another example of SATD declaration is the comment “# TODO handle attention history [^2] Accessed on: 2024-03-22, https://github.com/ufal/neuralmonkey/blob/master/neuralmonkey/attention/transformer_cross_layer.py within a code base that builds an attention model in the Neural Monkey project, since it indicates the missing implementation of handling model weights.
@@
The problem of TD in ML is exacerbated due to the “implicit” requirements of ML software products, which are grounded in data rather than specification documentations. As pointed out by Sculley et al. (Sculley), the problem of TD is pronounced in ML due to the stochastic nature of ML, as opposed to the deterministic nature of traditional (non-ML) software, leading to quality issues similar to traditional, non-ML software as well as new types of issues. For instance, an SATD in an ML model’s hyper-parameter configurations or data pre-processing will result in an inferior model with lower accuracy. Every incorrect decision made by such a model may potentially negatively affect various aspects such as the vendor’s reputation, user experience, or financial performance. Building on the insights of Sculley et al. (Sculley), this paper sets out to empirically validate whether the technical debt manifestations they describe in earlier work are indeed prevalent in real-world ML applications-and if so what are their characteristics. Empirically understanding the nature of SATD via code comments as recommended by Potdar and Shihbab allows us to explore direct insights into the acknowledged shortcomings in the implementation of ML systems. Since the nature of code changes performed in ML pipelines is different than that of the traditional non-ML software (bhatia2023towards), it is only reasonable to expect that SATD in ML code should appear distinct. Hence, we use code comments as a starting point, whereas other aspects of ML technical debt, such as debt in the underlying data, or model artifacts (perhaps embedded in project documentation, architectural design documentation, or project management systems) fall outside the scope of this current study. This delineation lays the groundwork for future research to explore such avenues.
@@
Despite recent efforts to understand SATD introduction and removal in deep learning frameworks, we still do not know much about the diffusion and evolution of technical debt in ML-based systems using such frameworks. Neither do we know the composition or the circumstances of SATD introduction or removal. While several researchers have investigated the detection, evolution, and impact of technical debt in traditional software systems (e.g., (Alfayez2020; Lim2012; liu2021exploratory; brown2010managing; wehaibi2016examining; letouzey2012managing; kruchten2012technical; kruchten2013technical; bavota2016large)), software systems powered by ML models (i.e., ML-based systems) have mostly remained out of scope.
@@
Figure 1. Example of SATD in a Natural Language Processing application, Mead-ML.
@@ 2.1. Studies on Software Engineering for Machine Learning
A systematic literature review by Nascimento et al. (nascimento2020software) highlighted the unique challenges in ML testing, ML software quality, and data management in ML systems. Moin et al. (moin2021data) proposed a novel approach to integrate Model-Driven Software Engineering and Model-Driven ML Engineering, with a focus on automated ML, aiming to assist software engineers without deep ML knowledge. Tantithamthavorn et al. (tantithamthavorn2018impact) emphasized the need for explainable ML in software engineering, making ML models more practical and actionable. Kästner and Kang (kastner2020teaching) discussed the importance of teaching software engineering skills to students with an ML background, focusing on realism, robust infrastructure, and ethical considerations. Singla et al. (singla2020story) analyzed the adoption of Agile methodology in machine learning projects, finding a higher number of exploratory tasks and difficulty in estimating task duration.
Within such a vast body of SE4ML research, our research expands on the “deployment maintenance and life-cycle management” area. This area is a ripe domain with prior research focusing on various aspects like code changes (e.g., Bhatia et al. (bhatia2023towards) examined the nature of code changes in open source ML software), ML asset management (Idowu et al. (idowu2022asset) focused on tracking, exploring, and retrieving ML assets), ML code smells (e.g., Jebnoun et al. (jebnoun2020scent) analyzed 10 code smells and their properties in ML software). In contrast, our study focuses on the SATD aspect and helps ML stakeholders understand the nature of TD and enable ML practitioners to gain awareness of TD in building ML applications.
@@ 2.2. Empirical Studies on SATD in Traditional (Non-ML) Software
In 1994, Cunningham (cunningham1992wycash) introduced the concept of technical debt as “quick and dirty” work in software design and implementation resulting in “not quite-right code”. However, TD is an abstract concept, as making tradeoffs between optimal software and meeting project deadlines is not quantitatively measurable. Hence, subsequently, Potdar and Shibab (potdar2014exploratory) extended the concept of TD to self-admitted TD (SATD), which was deemed to be a more visible measure of TD (nord2012search; kruchten2012technical; Alves2016). Along with introducing the concept of SATD, Potdar and Shibab examined SATDs in large open-source projects - Eclipse, Chromium, Apache HTTP, and ArgoUML. They found a high prevalence (within up to 31% of files) of SATD, and identified that more experienced developers add more SATD and are unable to pay off the technical development due to time pressure and code complexity.
@@
Building on the seminal work of Potdar and Shibab, multiple subsequent studies have identified SATD in traditional software. For instance, Maldonado et al. (maldonado2017empirical) studied the evolutionary aspects of traditional code and revealed that SATD lasts between 18 and 172 days in a system. Bavota and Russo (bavota2016large) performed a large-scale study of SATD across 159 software projects, contributing towards developing a taxonomy of SATD for traditional software. Wehaibi et al. (wehaibi2016examining) examined the impact of SATD on software quality, while Kashiwa et al. (kashiwa2022empirical) studied SATD in code reviews.
@@
However, none of the prior studies investigate SATD in ML-based systems. As identified by the previous subsection, the engineering practices and research-based findings of traditional software cannot be expected to hold as-is for ML-based systems (washizaki2019studying; ozkaya2020really; martinez2021software), and hence, our research is important to identify the characteristics of SATD in the currently prevalent ML-era of software development.
@@
Two studies bear similarities to our research. Firstly, akin to our study, the study by Liu et al. (liu2018satd) explores SATD in deep learning frameworks by investigating the introduction and removal of SATD in seven deep learning frameworks. However, contrary to the study of frameworks, our research identifies debt in ML systems, which use the former frameworks to build and integrate pipelines into ML applications. Our study provides insights for a broader spectrum of ML stakeholders who do not deal with ML mathematical or algorithmic implementations, but need to manage ML assets (like data, model) for development of ML applications. Secondly, the study by O’Brien et al. (obrien202223) mirrors our work closely. The authors check the different types of SATDs in ML software and check their distribution in different components of ML pipeline stages. While O’Brien et al.’s research is qualitative, we focus on both quantitative and qualitative aspects of SATD on a different dataset. Within the quantitative aspects, our work takes a genealogical approach to SATD, tracing the lifecycle from introduction to removal. Additionally, we put forth an exploratory model to understand the dynamics of long-lasting TD. Given our different dataset, our qualitative results complement the results of O’Brien et al. and provide a broader understanding of the prevalence and impact of debt across different datasets and domains.
@@ 2.3. Studies Presenting SATD Detection Tools
A few notable examples are the efforts made by Malando et al. (da2017using), who constructed an NLP-based SATD classification model and tested their tool across ten substantial open-source projects. Later, in 2017, Huang et al. (huang2018identifying) expanded the scope by developing a predictive SATD detection ML model that utilizes feature selection to provide the important words from the comment text used for SATD identification.
As an outcome of this research, the authors presented a composite SATD identification tool that combines multiple classifiers from different source code projects. Subsequently, Ren et al. (ren2019neural) utilized neural networks for SATD identification by leveraging the Convolutional Neural Network architecture to deliver state-of-the-art performance in detecting SATD comments. Later, in 2020, Faris et al. (de2020identifying) enhanced SATD classification models by using a contextualized vocabulary that considers the level of importance of different textual patterns along with the relationship between patterns and debt types.
@@
We employed the text mining-based SATD detector tool developed by Liu et al. (liu2018satd) in our study, which leverages natural language processing and machine learning to classify SATD comments. We selected Liu et al.’s SATD detector tool because of its frequent maintenance and demonstrated applicability in other authoritative SATD studies (e.g.,  (liu2021exploratory) and (Zampetti)).
@@ 3.1. RQ1: What is the prevalence of SATD in ML-based systems?
Sculley et al.’s seminal work (Sculley) highlighted the tendency of ML programs to incur technical debt, remarking that the strong dependence of ML models on data makes it difficult to enforce strict abstraction boundaries. When building ML-based systems, developers often experiment with different configurations in search of the optimal model; building, testing, and comparing the performance of different prototypes to identify the best configuration leading to the most efficient model (amershi2019software). This process can often result in multiple dead experimental code paths, increasing the cyclomatic complexity of the code and the risk of an unexpected behavior arising from some of the obsolete experimental code paths (Sculley). Another potential area where technical debt can accumulate quickly in ML code is in the model building or data processing configurations. Developers often have to handle a wide range of configuration options. For example, they often have to select among a variety of algorithm-specific learning settings, and–or from a large pool of pre-processing, post-processing, and verification techniques. In this research question, we aim to examine the extent to which ML software deployed today, contains technical debt, by comparing the prevalence of SATD in ML and non-ML code. While Sculley et al. (Sculley) elucidated additional ML challenges leading to accumulation of technical debt that are not present in non-ML software, empirical evaluations of said challenges still remain scarce. Hence, in this research question, we aim to examine the extent to which ML software deployed today, contains technical debt, by comparing the prevalence of SATD in ML and non-ML code.